# ZIP: Scalable Crowd Counting via Zero‑Inflated Poisson Modeling

### v2

@@ Conclusion
To assess the practical deployment potential of our lightweight models, we measure the average inference speed of each ZIP variant across four representative hardware platforms: Apple M1 Pro CPU (6 performance + 2 efficiency cores), Apple M1 Pro 14-core GPU, AMD Ryzen 9 5900X CPU (12-core), and NVIDIA RTX 3090 GPU. All experiments are conducted on the ShanghaiTech B dataset using full-precision inference at an input resolution of $1024\times 768$, which reflects real-world surveillance camera scenarios. The PyTorch version is 2.7.1. Table 6 reports the average frames per second (FPS) for each model. Our smallest variant, ZIP-P, achieves the highest throughput on all platforms, reaching 122.22 FPS on AMD 5900X and over 1,000 FPS on RTX 3090, making it ideal for resource-constrained or real-time applications. On the other end, ZIP-S delivers the best accuracy (MAE = 5.83) but at a higher computational cost, with CPU-side inference limited to under 5 FPS, making it suitable for accuracy-critical but latency-tolerant use cases. ZIP-N and ZIP-T offer compelling accuracy-efficiency trade-offs. For example, ZIP-T achieves 18.37 FPS on AMD 5900X and 249.73 FPS on RTX 3090, while maintaining a competitive MAE of 6.67. These results indicate that mid-sized variants can enable near real-time performance even on high-end CPUs without requiring dedicated GPUs. Overall, our ZIP family provides a scalable solution space, allowing deployment under varying hardware and latency constraints without compromising on count accuracy.
@@
ZIP (ours) 58.7 93.6 6.3 10.3 73.0 126.1 38.3 90.3
Improvement over EBC 5.9% 2.9% 10.0% 5.5% 5.4% 3.2% 3.2% 5.7%
Table 7: Comparison of ZIP with other frameworks on ShanghaiTech A & B, UCF-QNRF and NWPU-Crowd. All frameworks use the VGG19-based backbone. Results demonstrate that our framework ZIP significantly outperforms other methods, achieving the lowest MAE and RMSE on all four benchmarks. Our ZIP also consistently improves the EBC framework, at most by 10.0%. The best results are highlighted in bold font, and the second best results are underscored.
@@
Framework Comparison
We compare ZIP framework with several existing methods on the four benchmarks, shown in Table 7. Our approach sets a new state-of-the-art on ShanghaiTech B, UCF-QNRF and NWPU-Crowd. It also achieves the lowest MAE and the second lowest RMSE on ShanghaiTech A. Notably, when compared directly with its EBC baseline, our ZIP model demonstrates significant performance gains across all datasets, reducing the MAE by a margin of 3.2% to 10.0%. These improvements validate the efficacy of incorporating zero-inflated Poisson regression for crowd counting.
@@
Figure 7: Examples of annotation noise in NWPU-Crowd. Left column: input images with zoom-in views of the regions outlined in red. Center column: ground-truth density map with total count. Right column: density map with total count predicted by ZIP-B. Top row: mirror reflections labeled as real people. Bottom row: photo-realistic audience on a display screen left unlabeled.
Annotation Noise in NWPU-Crowd
Although NWPU-Crowd is currently the largest high-resolution benchmark, we observed non-negligible label noise. Fig. 7 presents two typical failure modes: over-labeled reflections (top row) and under-labeled screen display (bottom row).
@@
The bottom row depicts a conference hall is shown on a large display. The audience visible inside the screen looks photo-realistic, but none of these heads are annotated (GT = 221). The model nevertheless detects them, yielding a higher prediction (671). Qualitatively, the additional detections correspond to the people shown on the screen.
These examples underline two points: (i) reported errors can stem from annotation artifacts rather than model failure, and (ii) the structural-zero + ZIP formulation is robust enough to down-weight unrealistic cues (mirror images) while still detecting plausible but unlabeled heads (video feed). Future benchmarks would benefit from more consistent annotation protocols or confidence tags to flag uncertain regions.
@@
Despite its overall strong performance, ZIP-B still fails under certain visual ambiguities. Fig. 8 highlights two typical errors. The top row corresponds to false positive on man-shaped artifact: in an indoor museum scene, a post on the rope barrier strongly resembles a standing person in both shape and color, and as a result, ZIP-B misclassifies the post as a human. In the bottom row, rows of soldiers wear camouflage that blends with the background trucks. Although nearby soldiers are detected reliably, the model fails to pick up many small, low-contrast heads in the far field, leading to an under-count of -94. These failure cases suggest that incorporating additional sensing modalities, particularly infrared imagery, may help suppress person-like artifacts and recover low-contrast targets, thereby reducing both false positives and false negatives. Exploring how the ZIP formulation can be extended to multi-modal crowd counting will be an important direction for our future work.
@@
Blockwise Count Distribution
For each dataset’s training split, we form a pixel-level density map by placing 1 at each annotated head location and 0 elsewhere, without Gaussian smoothing being applied. We then slide an axis-aligned $B{\times}B$ window over the map with stride 1 pixel and record the integer count equal to the sum of densities in that window (equivalently, the number of annotated heads whose pixel coordinates fall inside the window). We repeat for $B\in\{8,16,32\}$ and aggregate counts over all windows from all images, converting to percentages. Fig. 9 illustrates the results. As the block size (window size) $B$ increases, zero percentages drop because each window integrates over a larger receptive field and is therefore more likely to include at least one head annotation. However, even at $B{=}32$ the distributions remain heavily zero-inflated. For instance, $>80\%$ zeros in UCF-QNRF and $>90\%$ in ShanghaiTech B and NWPU-Crowd. This indicates that annotation sparsity is a persistent regime rather than an artifact of using very fine windows. Note that increasing $B$ also coarsens supervision: spatial detail about where heads occur within a window is lost, making it harder for a crowd counting model to localize density peaks and learn fine-grained cues. In the extreme limit, $B$ equal to the entire image collapses supervision to a single global count label; the learning problem degenerates to (image-level) weakly supervised counting with no spatial guidance at all. Our ZIP formulation is motivated by operating in the practically useful regime of relatively small block size, where supervision is highly sparse yet still spatially informative, necessitating an explicit treatment of excess zeros.

### v3

@@ Conclusion
To assess the practical deployment potential of our lightweight models, we measure the average inference speed of each ZIP variant across four representative hardware platforms: Apple M1 Pro CPU (6 performance + 2 efficiency cores), Apple M1 Pro 14-core GPU, AMD Ryzen 9 5900X CPU (12-core), and NVIDIA RTX 3090 GPU. All experiments are conducted on the ShanghaiTech B dataset using full-precision inference at an input resolution of 1024×7681024$\times$ 7681024 × 768, which reflects real-world surveillance camera scenarios. Inference speed is computed as the reciprocal of the average per-image inference time, where the time is measured individually for each test image. The PyTorch version is 2.7.1. Table 6 reports the average frames per second (FPS) for each model. Our smallest variant, ZIP-P, achieves the highest throughput on all platforms, reaching 122.22 FPS on AMD 5900X and over 1,000 FPS on RTX 3090, making it ideal for resource-constrained or real-time applications. On the other end, ZIP-S delivers the best accuracy (MAE = 5.83) but at a higher computational cost, with CPU-side inference limited to under 5 FPS, making it suitable for accuracy-critical but latency-tolerant use cases. ZIP-N and ZIP-T offer compelling accuracy-efficiency trade-offs. For example, ZIP-T achieves 18.37 FPS on AMD 5900X and 249.73 FPS on RTX 3090, while maintaining a competitive MAE of 6.67. These results indicate that mid-sized variants can enable near real-time performance even on high-end CPUs without requiring dedicated GPUs. Overall, our ZIP family provides a scalable solution space, allowing deployment under varying hardware and latency constraints without compromising on count accuracy.
Datasets and Augmentation
Datasets. ShanghaiTech A & B (Zhang et al. 2016), UCF-QNRF (Idrees et al. 2018), and NWPU-Crowd (Wang et al. 2020c) are four widely used crowd counting benchmarks. ShanghaiTech A contains 300 training images and 182 test images, with highly congested scenes and an average of 501 people per image. ShanghaiTech B, the only dataset collected from surveillance viewpoints, includes 400 training and 316 test images, featuring relatively sparse crowds. UCF-QNRF comprises 1,201 training images and 334 test images with extremely dense crowds, averaging 815 people per image. NWPU-Crowd is currently the largest high-resolution crowd counting dataset, containing 3,109 training images, 500 validation images, and 1,500 test images, with an average count of 418. The test set annotations of NWPU-Crowd are not publicly released. To obtain evaluation results on this split, predicted counts must be submitted to the official evaluation server. This protocol helps maintain a fair comparison between methods by preventing result tweaking using test labels.
We apply a combination of data augmentation techniques during training, largely following the EBC framework (Ma, Sanchez, and Guha 2024). The transformation pipeline includes the following operations:
[…]
  • RandomHorizontalFlip: The image is flipped horizontally with a probability of 0.5.
  • ColorJitter: Brightness, contrast, and saturation are independently perturbed within specified ranges to simulate varying lighting and color conditions. Hue adjustment is disabled (set to 0.0). We use the implementation from torchvision to achieve this augmentation.
  • PepperSaltNoise: A small fraction of pixels are randomly replaced with either black (salt, value 0) or white (pepper, value 1), introducing local binary noise. The noise levels are controlled by two parameters: saltiness (ratio of salt pixels) and spiciness (ratio of pepper pixels). These two parameters are fixed to 0.001 across all datasets.
We use different values for training resolution, block sizes and augmentation ranges, and the details are presented in Table 7.
Method Luminance-Level Scene-Level
L1 L2 L3 S1 S2 S3 S4
CrowdHat (Wu and Yang 2023) 19.13 19.63 13.04 19.36 18.62 14.64 21.69
STEERER (Han et al. 2023) 14.37 14.26 9.90 16.67 11.28 12.22 23.66
APGCC (Chen et al. 2024) 31.57 22.07 17.28 36.10 14.72 14.13 29.33
CLIP-EBC (Ma, Sanchez, and Guha 2024) 16.47 15.55 11.88 20.69 11.81 12.10 22.13
ZIP-B (ours) 14.49 11.23 8.26 11.13 9.97 11.61 21.57
Table 8: NAE (%) comparison of ZIP-B with the latest methods on the NWPU-Crowd test set under different luminance levels and scene densities. The best result within each subgroup is shown in bold, and the second best is underlined. ZIP-B achieves the lowest NAE across most subgroups, demonstrating robustness to lighting and crowd scale variations.
Bin Configurations
To accommodate varying count distributions across datasets, the blockwise ground-truth counts were quantized into dataset-specific integer bins. Each bin is defined as a closed interval $[a,b]$ corresponding to a discrete count range, and its associated bin center is used to estimate the expected value within the ZIP framework. The following configurations were adopted for each dataset:
  • ShanghaiTech A (block size = 16): The counts were discretized into 14 bins: 11 singleton bins for counts from 0 to 10, two intermediate bins [11,12][11,12][ 11 , 12 ] and [13,14][13,14][ 13 , 14 ], and one final bin [15,)[15$\infty$)[ 15 , ∞ ) to cover the long tail of larger counts. The corresponding bin centers were: 0.0, 1.0, 2.0, 3.0, 4.0, 5.0, 6.0, 7.0, 8.0, 9.0, 10.0, 11.38, 13.38, and 16.26, where the final three values were computed as the empirical means within each multi-count interval.
  • ShanghaiTech B (block size = 16): A more compact binning scheme was used, reflecting the relatively sparse nature of the dataset. Ten bins were defined: eight singleton bins for counts 0 through 8 and one open-ended bin [9,)[9$\infty$)[ 9 , ∞ ). The bin centers were: 0.0, 1.0, 2.0, 3.0, 4.0, 5.0, 6.0, 7.0, 8.0, 10.16.
  • UCF-QNRF (block size = 32): To model the highly dense and skewed count distribution, a fine-grained binning strategy was adopted. The counts were divided into 21 bins: 11 singleton bins (0–10), followed by progressively wider intervals up to [34,)[34$\infty$)[ 34 , ∞ ). The corresponding bin centers included both integers and empirically estimated means for the aggregated bins: 0.0, 1.0, $\cdots$, 10.0, 11.43, 13.43, 15.44, 17.44, 19.43, 21.83, 24.85, 27.87, 31.24, 38.86.
  • NWPU-Crowd (block size = 16): A simplified 11-bin scheme was employed, with singleton bins from 0 to 9 and one catch-all bin [10,)[10$\infty$)[ 10 , ∞ ) to handle outliers. The bin centers were: 0.0, 1.0, $\cdots$, 9.0, 12.16, where the last center reflects the average count in the aggregated tail.
Luminance-Level and Scene-Level Comparison
To further evaluate the robustness and generalizability of ZIP-B, we conduct a detailed analysis on the NWPU-Crowd test set, stratifying the results by both luminance levels and scene density levels. Luminance was measured using the average pixel intensity in the Y channel of the YUV color space, and divided into three groups: L1 [0,0.25][0,0.25][ 0 , 0.25 ], L2 (0.25,0.5](0.25,0.5]( 0.25 , 0.5 ], and L3 (0.5,0.75](0.5,0.75]( 0.5 , 0.75 ]. Scene-level difficulty was assessed based on the ground-truth number of people per image, grouped into four intervals: S1 (0,100](0,100]( 0 , 100 ], S2 (100,500](100,500]( 100 , 500 ], S3 (500,5000](500,5000]( 500 , 5000 ] and S4 (>5000>5000> 5000). Performance was measured using Normalized Absolute Error (NAE), reported in percentage.
As shown in Table 8, ZIP-B consistently outperforms prior methods across six out of seven subgroups, achieving the lowest NAE in L2 (11.23%), L3 (8.26%), and all four scene levels (S1–S4), showing that ZIP-B is not only accurate on average, but also robust across varying lighting conditions and crowd densities. In particular, it shows notable gains under sparse scenes (S1), demonstrating the effectiveness of the zero-inflated Poisson modeling.
Training Size Block Size
16 32
MAE RMSE MAE RMSE
448 78.52 130.24 75.73 127.17
672 75.06 131.62 73.02 126.11
896 77.92 133.65 75.69 130.30
Table 9: Impact of training resolution and block size on UCF-QNRF with the VGG19-based backbone. Using a larger block size (32) consistently improves MAE and RMSE across resolutions. The best overall performance is achieved with a resolution of 672 and block size 32, highlighting the benefit of coarser spatial supervision for high-resolution dense crowd scenes.
Effect of Training Resolution and Block Size on UCF-QNRF
We study the impact of varying the training input resolution and block size on model performance using the UCF-QNRF dataset and the VGG19-based encoder-decoder structure. The results are presented in Table 9.
Across all training resolutions, using a larger block size of 32 consistently yields better performance than a block size of 16 in terms of both MAE and RMSE. For instance, at a resolution of 672, the configuration with block size 32 achieves the best overall performance, with an MAE of 73.02 and a RMSE of 126.11. This suggests that for UCF-QNRF–a dataset with high-resolution images and dense crowd scenes–larger blocks better capture spatially aggregated count patterns, possibly providing more stable supervision.
While increasing the training resolution generally improves RMSE slightly, the gains saturate or diminish beyond 672. This might be attributed to the sampling behavior of RandomResizedCrop and how it interacts with image and block sizes. At lower resolutions (e.g., 448), cropped regions often lack sufficient spatial coverage to include high-density areas, resulting in a limited number of blocks with large counts. This weakens supervision on the tail of the count distribution. On the other hand, at very high resolutions (e.g., 896), the upper end of the cropping scale becomes ineffective, as attempting to crop regions larger than the image leads to upscaling followed by near-trivial cropping and resizing. This process fails to increase the sample sizes of large blockwise counts and may introduce resampling artifacts, which together limit the benefits of increased resolution. These observations explain why intermediate resolutions (e.g., 672) paired with a coarser block size (e.g., 32) achieve the best overall results on UCF-QNRF.
@@
ZIP (ours) 58.7 93.6 6.0 9.9 73.0 126.1 38.3 90.3
Improvement over EBC 5.9% 2.9% 14.2% 9.1% 5.4% 3.2% 3.2% 5.7%
Table 10: Overall comparison of ZIP with other frameworks on ShanghaiTech A & B, UCF-QNRF and NWPU-Crowd. All frameworks use the VGG19-based backbone. Results demonstrate that our framework ZIP significantly outperforms other methods, achieving the lowest MAE and RMSE on all four benchmarks. Our ZIP also consistently improves the EBC framework, at most by 14.2%. The best results are highlighted in bold font, and the second best results are underscored.
@@
Robustness to Random Seed Initialization
To assess the robustness of the proposed ZIP-B model with respect to random initialization, an additional experiment was conducted by training ZIP-B on the NWPU-Crowd validation split using five different random seeds: 1, 42, 1234, 3407, and 12345. This dataset was selected due to its large size (500 instances), which offers a more statistically stable evaluation compared to other benchmarks. The model achieved a mean MAE of 29.00 with a standard deviation of 0.83, indicating low sensitivity to random initialization. For comparison, the result reported in Table 3 for ZIP-B on NWPU-Crowd (val) is a single-run MAE of 28.2. The fact that this value falls within one standard deviation of the mean observed across five seeds confirms that the previously reported performance is representative and not an outlier due to a favorable random initialization. These results underscore the stability and reliability of the ZIP framework under common sources of stochasticity during training.
Overall Comparison of Frameworks
We compare ZIP with several existing frameworks on the four benchmarks, shown in Table 10. All methods are based on the same VGG19 backbone. Our approach sets a new state-of-the-art on ShanghaiTech A & B, UCF-QNRF and NWPU-Crowd. Notably, when compared directly with its EBC baseline, our ZIP model demonstrates significant performance gains across all datasets, reducing the MAE by a margin of 3.2% to 14.2%. These improvements validate the efficacy of zero-inflated Poisson regression for counting.
@@
Figure 7: Examples of annotation noise in NWPU-Crowd. Left column: input images with zoom-in views of the regions outlined in red. Center column: ground-truth density map with total count. Right column: density map with total count predicted by ZIP-B. Top row: mirror reflections labeled as real people. Middle row: photo-realistic audience on a display screen left unlabeled. Bottom row: snow-covered tree branches mislabeled as a dense crowd.
Annotation Noise
Although NWPU-Crowd is currently the largest high-resolution benchmark, we observed non-negligible label noise. Fig. 7 presents three typical failure modes on the validation split of this dataset: over-labeled reflections (top row), under-labeled screen display (middle row), and complex natural textures being mislabeled as crowds (bottom row).
@@
The middle row depicts a conference hall is shown on a large display. The audience visible inside the screen looks photo-realistic, but none of these heads are annotated (GT = 221). The model nevertheless detects them, yielding a higher prediction (671). Qualitatively, the additional detections correspond to the people shown on the screen.
In the bottom row, a snowy park scene is shown where the red box highlights complex textures on snow-covered tree branches. These natural textures have been mistakenly annotated as a dense crowd in the background, leading to a grossly overestimated ground truth count of 993. Our model is more robust to this type of noise, largely ignoring the mislabeled texture and predicting a more plausible count of 448.
These examples underline two points: (i) reported errors can stem from annotation artifacts rather than model failure, and (ii) the structural-zero + ZIP formulation is robust enough to down-weight unrealistic cues while still detecting plausible but unlabeled heads. Future benchmarks would benefit from more consistent annotation protocols, confidence tags to flag uncertain regions, and a multi-annotator verification process to improve label quality and consistency.
@@
Despite its overall strong performance, ZIP-B still fails under certain visual ambiguities. Fig. 8 highlights two typical errors. The top row corresponds to false positive on man-shaped artifact: in an indoor museum scene, a post along the rope barrier closely resembles a standing person in both shape and color, and as a result, ZIP-B misclassifies the post as a human. In the bottom row, rows of soldiers wear camouflage that blends with the background trucks. Although nearby soldiers are detected reliably, the model fails to pick up many small, low-contrast heads in the far field, leading to an under-count of -94. These failure cases suggest that incorporating additional sensing modalities, particularly infrared imagery, may help suppress person-like artifacts and recover low-contrast targets, thereby reducing both false positives and false negatives. Exploring how the ZIP formulation can be extended to multi-modal crowd counting will be an important direction for our future work.
Blockwise Count Distribution
[…]
@@
(a) Input Image.
(b) Structural Zero Map $\boldsymbol{\pi}$.
(c) Lambda Map $\boldsymbol{\lambda}$.
(d) Sampling Zero Map.
(e) Complete Zero Map.
(f) Predicted Density Map.
Figure 10: Visualization of the interpretability of our ZIP framework. This figure decomposes ZIP-B’s prediction on a test image from ShanghaiTech B into its key components, illustrating how the model separates structurally irrelevant regions from informative ones. (a) Input Image: A representative test image from the ShanghaiTech B dataset. (b) Structural Zero Map $\boldsymbol{\pi}$: Shows the model’s estimated probability that a block is structurally empty (e.g., background, non-head regions). Red regions indicate high structural-zero probability, while blue/cyan spots highlight likely head-center regions, effectively masking irrelevant areas like pavement. (c) Lambda Map $\boldsymbol{\lambda}$: The predicted Poisson rate map, representing expected counts per block. (d) Sampling Zero Map: Visualizes the probability of zero counts due to sampling error (i.e., head-center blocks receive zero counts due to the point-based labels). Computed as $(1-\boldsymbol{\pi})\otimes\exp(-\boldsymbol{\lambda})$, it peaks at sparser head-center regions. (e) Complete Zero Map: The total probability of a zero count from any source, combining both structural and sampling zeros. (f) Predicted Density Map $\boldsymbol{Y}^{*}$: The final predicted density map, computed as $(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda}$, integrating both spatial suppression and expected counts. Together, these visualizations show how ZIP disentangles where people are located from how densely they are present.
Model Interpretation
A key advantage of the proposed Zero-Inflated Poisson (ZIP) framework is its inherent interpretability. By decomposing the counting process into distinct probabilistic components, ZIP enables us to diagnose the model’s behavior by separately analyzing where it focuses and how much it counts, as illustrated in Fig. 10.
The structural zero map ($\boldsymbol{\pi}$) functions as a learned attention or segmentation mechanism. This branch estimates the probability that a given block is structurally irrelevant to crowd counting (e.g., background, buildings, torso, or peripheral head parts) and should be suppressed. As shown in Fig. 10(b), the model assigns high structural-zero probability (deep red) to background areas like pavement, while assigning low probability (blue/cyan) to compact regions containing head centers. Note that this spatial disentanglement is learned without any segmentation supervision, showing that ZIP can implicitly separate relevant and irrelevant regions using only point-level annotations.
In parallel, the lambda map $\boldsymbol{\lambda}$ captures the expected count per block. As visualized in Fig. 10(c), high values appear in the upper part of the image where the crowd is denser due to perspective.
The sampling zero map (Fig. 10(d)) models the probability that a block receives a zero count due to sampling effects (i.e., head-center blocks receive a zero counts because of the point-based annotation). It is computed as $(1-\boldsymbol{\pi})\otimes\exp(-\lambda)$, and is highest where the model is confident a person is present (low $\boldsymbol{\pi}$), but the predicted count is modest (moderate $\lambda$).
The complete zero probability map (Fig. 10(e)) represents the overall likelihood that a block receives a zero count from any source (either structural or sampling). It is defined as defined as $\boldsymbol{\pi}+(1-\boldsymbol{\pi})\otimes\exp(-\lambda)$. This map is often dominated by the structural zero component, as the background makes up the majority of the image. It can be viewed as an inverse attention map, where non-red regions highlight areas the model expects to contain people.
Finally, the predicted density map $\boldsymbol{Y}^{*}$ (Fig. 10(f)) is computed as $(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda}$, combining spatial suppression and count estimation. This yields a clean and well-localized density prediction, with high activations concentrated near annotated head centers.
Together, these components demonstrate how ZIP not only improves counting accuracy, but also offers fine-grained interpretability by isolating both spatial relevance and count magnitude across the image.